OpenStudy (ammarah):

Find the vertex for the quadratic function. u(x)= 0.5x^2-2x+5

OpenStudy (ammarah):

@ganeshie8

OpenStudy (anonymous):

first coordinate is \(-\frac{b}{2a}\)

OpenStudy (ammarah):

1? i dont get how to set it up

OpenStudy (anonymous):

as you have u(x)= 0.5x^2-2x+5 here a = 0.5, b = -2 and c = 5 to find the first coordinate of vertex we have x = -b/(2a) find x and put it in u(x)= 0.5x^2-2x+5 to find u(x) = y thus you will get both coordinates (x,y) =(........,.........)

OpenStudy (ammarah):

no no i mean how to put it in vertex "form"

OpenStudy (ammarah):

dont u first subtract 5 so it would be:

OpenStudy (mayankdevnani):

@Ammarah To find VERTEX,we have its co=ordinates as :- \[\large \bf Vertex(\frac{-b}{2a},\frac{-D}{4a})\]

OpenStudy (triciaal):

a(x-b/2a)^2 + k =

OpenStudy (mayankdevnani):

first write in standard form(ax^2+by+c=0) and then use that formula

OpenStudy (ammarah):

it is in tat form/!!!

OpenStudy (mayankdevnani):

correct and then use that formula

OpenStudy (ammarah):

yes my question is how to write the vertex form all u guys r doing is teling me which formula to use@

OpenStudy (anonymous):

\(-\frac{b}{2a}=2\) if i am not mistaken

OpenStudy (anonymous):

is that correct?

OpenStudy (mayankdevnani):

@Ammarah use that formula(in attachment)

OpenStudy (anonymous):

i wouldn't

OpenStudy (mayankdevnani):

you are correct.. -b/2a=2

OpenStudy (anonymous):

we get \[-\frac{b}{2a}=2,u(2)=3\] so the vertex is \((2,3)\) and therefore the vertex form is \[.5(x-2)^2+3\]halas

OpenStudy (ammarah):

umm -2/2^2 is 3?

OpenStudy (anonymous):

i must have lost you \[-\frac{b}{2a}=-\frac{-2}{2\times .5}=\frac{2}{1}=2\]

OpenStudy (mayankdevnani):

@satellite73 he want to find the co-ordinates of VERTEX

OpenStudy (mayankdevnani):

not its form

OpenStudy (anonymous):

\[.5x^2-2x+5\] \[ax^2+bx+c\] \[a=.5,b=-2,c=5\]

OpenStudy (ammarah):

ok please show step by step

OpenStudy (anonymous):

\[-\frac{b}{2a}=...=2\]and \[.5\times 2^2-2\times 2+5=3\]

OpenStudy (ammarah):

u lost

OpenStudy (ammarah):

u lost me at the times part

OpenStudy (mayankdevnani):

@Ammarah my solution is correct and appropriate

OpenStudy (mayankdevnani):

your question wants to say that you have to find the vertex(its co-ordinates)

OpenStudy (ammarah):

−b2a=2,u(2)=3 so the vertex is (2,3) and therefore the vertex form is .5(x−2)2+3 I dont get how you just skipped to the .5 (x-2) part please show how u got there

OpenStudy (mayankdevnani):

first ,find its x-coordinate

OpenStudy (ammarah):

just show me............i will uderstand if u just show me the steps

OpenStudy (mayankdevnani):

Vertex's x coordinate=-b/2a where b=-2 and a=0.5 after simplify,we get \[\large \bf \frac{-b}{2a}=-\frac{-2}{0.5}=2\]

OpenStudy (mayankdevnani):

understood till here ?